OpenStudy (anonymous):

Use the half-angle formulas to come up with an exact expression for each formula value below. You do not have to simplify your answers. a.) cos(pi/4) b.)cos(pi/8) c.)cos(pi/16)

OpenStudy (anonymous):

How would you find this to simplify each one?

OpenStudy (tkhunny):

?? Why would you use a half-angle identity to find \(\cos(\pi/4) = \dfrac{\sqrt{2}}{2}\)?

OpenStudy (anonymous):

Thats the question.... I'm trying to figure that out as well

OpenStudy (anonymous):

I think the answer is supposed to be in degrees that is why

OpenStudy (tkhunny):

Well, I vote for skipping that one and starting with \(\pi/8\) Since everything is in Quadrant I, let's just dive in. We have the Double Angle Formula \(\cos(2x) = 2\cos^{2}(x) - 1\). Demoting to Half Angle, we have:\(\cos(x) = 2\cos^{2}(x/2) - 1\). Solving for the half angle gives: \(\cos(x/2) = \sqrt{\dfrac{\cos(x) + 1}{2}}\). Using \(x = \pi/4\), we have: \(\cos(\pi/8) = \sqrt{\dfrac{\cos(\pi/4) + 1}{2}}\). Wade through all that and then show me \(\cos(\pi/16)\)

OpenStudy (anonymous):

Wouldn't it be the same thing but you would have the pi/16 numbers involved? because the decimal answer is .98

OpenStudy (tkhunny):

No decimals. "with an exact expression for " <== It's in the problem statement. Put your calculator away!

OpenStudy (anonymous):

Ok, but it says you don't have to simplify.. So I was right about the pi/16

OpenStudy (tkhunny):

Not if you think it should look like "0.98". Maybe you could just rewrite my last expression with pi/16 on the outside and pi/8 on the inside? It's not perfectly clear what it means not to simplify.

OpenStudy (anonymous):

Alright, Thank you!

OpenStudy (anonymous):

@Ivlady you ok?